# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ III-G Racing
Autonomous racing represents a distinctive facet of autonomous driving research, demanding vehicles to leverage intricate algorithms, a suite of sensors, and cutting-edge control systems. This specialized form of driving involves navigating tracks, engaging in competitive races, and overcoming dynamic racing challenges. Unlike standard driving tasks, racing pushes vehicles to attain high velocities, extensively testing the dynamic limits of these automated systems. When racing at high speeds, vehicles must swiftly detect other vehicles or obstacles, demanding rapid reaction times. Additionally, they must accurately localize their position concerning the track and strategize dynamic trajectories to optimize performance [211]. While extensive research has been dedicated to full-size racing competitions like Roborace [212], our primary focus centers on small-scale car racing events such as AutoRally [9], F1TENTH [12], Donkeycar [11], ORCA [15] and others.
III-G1 Modular system
In this section, we will explore the racing task using the modular system, which similar to other tasks, primarily comprises two essential modules: the perception and control modules. The perception module plays a critical role in racing tasks as it assists the vehicle in determining its position on the racetrack. In the racing context, swift and accurate reactions are essential to fulfill subsequent control tasks. Thus, a racing-specific perception module must balance speed and precision to enable quick and accurate responses during racing scenarios. For this task, the most commonly used perception module is the fusion with GPS and IMU as discussed in Section III-A. However, when states or sensor measurements that are not approximately linear in the measurement time frame, [9] and [51] use factor graphs combined with incremental smoothing and mapping 2 (iSAM2) [213], to fuse GPS and IMU measurements and output smoothing estimation. For the issues with non-linear dynamics and non-linear measurement models, [55, 54, 53] use the particle filter to fuse the camera images with IMU, and GPS. Additionally, a single camera can serve as a perception sensor for racing tasks. For example, in [56], CNN is used with a single monocular camera to predict the cost map of the track in front of the vehicle which is directly useable for online trajectory optimization with MPC. Similarly, [214] uses a top-down lane cost map CNN and the YOLOv2 CNN to extract feature-input values and a two-point visual driver control model (TPVDCM) as the controller to control the vehicle.
Figure 10: In [56], the cost map of the track in front of the vehicle is predicted with CNN.
@@ III-G1 Modular system
Same as the perception module, the control module also has a high demand for timely reactions, often surpassing the importance of the perception module. The most commonly used control module in racing is MPC, which is a control strategy that utilizes a dynamic model of the system to predict its future behavior and make control decisions based on optimization criteria. In [15], the authors use onboard IMU sensors and a global localization system to acquire the states of the car within the ORAC platform. Subsequently, a path planner and a Nonlinear MPC (NMPC) are utilized to guide the racing process. Later, in [51], a sampling-based MPC algorithm called the Model Predictive Path Integral Control algorithm (MPPI) is introduced for Autorally. This algorithm presents a novel derivation of path integral control, offering an explicit formula for controls across the entire time horizon. A notable attribute of MPPI is its capability to generate entirely new behaviors dynamically, enabling the controller to drive the vehicle to its operational limits. Then, a few improved versions are proposed, such as the robust sampling-based MPC framework based on a combination of model predictive path integral control and nonlinear Tube-MPC [215] and best response model predictive control based on a combination of the game-theoretic notion of iterated best response, and an information-theoretic model predictive control algorithm [216]. When the dynamic model lacks perfect accuracy, it introduces potential disparities between the predicted and actual behavior of the system. To address this challenge, Learning Model Predictive Control (LMPC) integrates learning algorithms with MPC, aiming to refine control strategies in the face of uncertain or changing dynamics. Compared to traditional MPC, LMPC continually enhances its performance, especially in scenarios where complete knowledge of system dynamics is unavailable or subject to variation. The learning mechanism algorithms such as RL, Gaussian processes, neural networks, or other adaptive learning methods in LMPC continuously update the predictive model based on collected data and feedback from the system, to achieve better performance, adaptability to changing system dynamics, and robustness in scenarios where precise models might be unavailable or incomplete. [53] establishes the connection between MPC and online learning, and proposes a new algorithm based on dynamic mirror descent (DMD) and MPC (DMD-MPC), which provides a fresh perspective on previous heuristics used in MPC. In [217], historical data is used to construct secure sets and approximate the value function, facilitating LMPC to learn and improve from past experiences within BARC platform. Later, [218] applies Gaussian processes to correct the model mismatch, then uses MPC for tracking pre-computed racing lines using this corrected model for an F1TENTH car. For more responsive control, [110] uses Linear Parameter Varying (LPV) theory to model the dynamics of the vehicle and combine it with MPC (LPV-MPC), which can be computed online with reduced computational cost. To swiftly identify unsafe conditions, [54] propose a Perceptual Attention-based Predictive Control (PAPC) algorithm, where MPC is first used to learn how to place attention on relevant areas of the visual input and ROI, and output control actions as well as estimates of epistemic and aleatoric uncertainty in the attention-aware visual input of an Autorally car. [219] introduces a local, linear, data-driven learning method for error dynamics within the LMPC framework. This approach exhibits increased robustness against parameter variations and limited data compared to prior LMPC implementations. Pure pursuit is another control algorithm used in racing tasks [220]. A key limitation in model-based control methods is that they require localization during the race, where the motion capture system is sometimes used [221, 222], but it is still limited by requiring a map of the track or a set of cameras and the hardware requirement for real-time localization.
III-G2 End-to-end system
Figure 11: End-to-end IL system used in [91] for an Autorally racing car.
In addressing the challenges within modular systems and enhancing racing performance, various studies have proposed an end-to-end system that incorporates methodologies such as IL and RL [223]. The authors in [91] introduce an end-to-end IL system depicted in Fig. 11, where a learner network needs to imitate an expert. The expert fuses GPS and IMU for state estimation and uses an MPC as controller, the learner uses a DNN as control policy to map raw, high-dimensional observations to continuous steering and throttle commands. [224] presents a deep imitative RL framework for end-to-end racing with camera input, where IL is used to initialize the policy, and model-based RL is used for further refinement by interacting with an uncertainty-aware world model. To improve the stability of the racing system, [225] designs a residual control system, in which multiple Bayesian Neural Networks (BNNs) are trained with two camera inputs and GPS measurements to control an Autorally car in an end-to-end fashion. Incorporating LiDAR technology into end-to-end systems has also garnered attention. A model-based RL approach is employed in [226] using raw LiDAR input on the F1TENTH car to navigate racetracks effectively. Subsequently, in [227], LiDAR data is used as input for DQN to control the F1TENTH car. This study also provides a comparative analysis between neural network methodologies for processing LiDAR data and evaluates two Sim2Real approaches. Although RL can be used for racing tasks, the well-known low sample efficiency issue still holds back further development. For this issue, [228] propose an efficient residual policy learning method with the raw observation of LiDAR and IMU, in which first a controller based on the modified artificial potential field (MAPF) is used to generate policies, then DRL algorithms are used to generate a residual policy as a supplement to obtain the optimal policy with increased efficiency. Similarly, [229] presents a trajectory-aided learning (TAL) method that trains DRL with raw LiDAR input by incorporating the optimal trajectory into the learning formulation. [230] also presents a residual vehicle controller that learns to amend a traditional controller with a similar idea. To tackle the safety issues that usually occur in RL training, [231] uses a Viability Theory-based supervisor to recursively feasible vehicle safety during the training. Furthermore, to improve the robustness of RL, [232] first train a teacher model that overfits the training track, moving along a near-optimal path, then use this model to teach a student PPO model the correct actions along with randomization.
III-G3 Drifting
Autonomous drifting can be seen only as a facet of racing, but also as an individual task within autonomous vehicle control. This maneuver is characterized by intentionally inducing oversteer, resulting in the rear tires losing traction and enabling the vehicle to elegantly slide or ”drift” through a turn. Autonomous drifting requires advanced control systems and algorithms that can precisely manage the speed, steering, throttle, and braking to maintain control while deliberately sliding through a corner. Few studies work on the drifting problems for small-scale cars, the key point is the awareness of the kinematic model and dynamic of the vehicle and the accurate estimation of the state such as position, and steering, to control the car appropriately. For instance, [8] uses a dynamic bicycle model for a BARC car, which has a single rigid body model with lumped rear wheels and front wheels. Then EKF is utilized to fuse the sensor measurement from IMU and encoders with the dynamic bicycle model and Pajecka tire model [233]. Moreover, optical flow is used to estimate the velocity, and a Linear-quadratic regulator (LQR) controller with designed equilibrium drifting points is used to control the vehicle. Later in [234], EKF is also applied for state estimation but with a six-state bicycle model with linear front and rear-wheel tire forces. In most cases, an accurate vehicle dynamics model is often elusive. Hence, [235] proposes an approach to tackle the drifting park problem with BARC cars by segmenting it into distinct phases: the normal driving regime and the sliding regime. During the normal driving phase, a nonlinear MPC operates with a predefined kinematic model. As the vehicle transitions into the sliding phase, reliance on this model diminishes. To navigate this shift, a feedforward-feedback controller takes charge, orchestrating safer maneuvers adeptly under sliding conditions. Another challenge in previous studies pertains to the singularities embedded within popular tire models. Specifically, these models exhibit singularity issues, notably with tire slip angles becoming singular at lower vehicle speeds due to the presence of the vehicle velocity term in the denominator. In response, [236] introduces a fused kinematic-dynamic bicycle model accompanied by a nonlinear MPC system for a RACECAR. This pioneering approach harmonizes the planning and execution of dynamic vehicle maneuvers within a unified framework, optimizing the entire process cohesively.

### v2

@@ IV-C5 Racing
Autonomous racing, distinct from traditional autonomous driving, emphasizes high-speed navigation, rapid reaction, and dynamic trajectory planning. Unlike standard driving tasks, racing pushes vehicles to attain high velocities, extensively testing the dynamic limits of these automated systems. When racing at high speeds, vehicles must swiftly detect other vehicles or obstacles, demanding rapid reaction times. Additionally, they must accurately localize their position concerning the track and strategize dynamic trajectories to optimize performance [246]. While extensive research has been dedicated to full-size racing competitions like Roborace, our primary focus centers on small-scale car racing events such as AutoRally [11], F1TENTH [12], Donkeycar, ORCA [20] and others.
For racing tasks, the control module has a high demand for timely reactions, often surpassing the importance of the perception module. The most commonly used control module in racing is MPC, which is a control strategy that utilizes a dynamic model of the system to predict its future behavior and make control decisions based on optimization criteria. In [20], a path planner and a Nonlinear MPC (NMPC) are utilized to guide the racing process within the ORAC platform. Later, in [50], a sampling-based MPC algorithm called the Model Predictive Path Integral Control algorithm (MPPI) is introduced for Autorally. This algorithm presents a novel derivation of path integral control, offering an explicit formula for controls across the entire time horizon. A notable attribute of MPPI is its capability to generate entirely new behaviors dynamically, enabling the controller to drive the vehicle to its operational limits. Then, a few improved versions are proposed, such as the robust sampling-based MPC framework based on a combination of model predictive path integral control and nonlinear Tube-MPC [215] and best response model predictive control based on a combination of the game-theoretic notion of iterated best response, and an information-theoretic model predictive control algorithm [247]. Despite its effectiveness in dynamic trajectory planning and adaptability to high-speed scenarios, MPC has notable drawbacks. It depends heavily on accurate dynamic models, which, if imperfect, can lead to disparities between predicted and actual behavior. Additionally, the high computational demand poses challenges for real-time implementation.
@@
To address this challenge, Learning Model Predictive Control (LMPC) integrates learning algorithms with MPC, aiming to refine control strategies in the face of uncertain or changing dynamics. Compared to traditional MPC, LMPC continually enhances its performance, especially in scenarios where complete knowledge of system dynamics is unavailable or subject to variation. The learning mechanism algorithms such as RL, Gaussian processes, neural networks, or other adaptive learning methods in LMPC continuously update the predictive model based on collected data and feedback from the system, to achieve better performance, adaptability to changing system dynamics, and robustness in scenarios where precise models might be unavailable or incomplete. [52] establishes the connection between MPC and online learning, and proposes a new algorithm based on dynamic mirror descent (DMD) and MPC (DMD-MPC), which provides a fresh perspective on previous heuristics used in MPC. In [216], historical data is used to construct secure sets and approximate the value function, facilitating LMPC to learn and improve from past experiences within BARC platform. Later, [214] applies Gaussian processes to correct the model mismatch, then uses MPC for tracking pre-computed racing lines using this corrected model for an F1TENTH car. For more responsive control, [121] uses Linear Parameter Varying (LPV) theory to model the dynamics of the vehicle and combine it with MPC (LPV-MPC), which can be computed online with reduced computational cost. To swiftly identify unsafe conditions, [53] propose a Perceptual Attention-based Predictive Control (PAPC) algorithm, where MPC is first used to learn how to place attention on relevant areas of the visual input and ROI, and output control actions as well as estimates of epistemic and aleatoric uncertainty in the attention-aware visual input of an Autorally car. [146] introduces a local, linear, data-driven learning method for error dynamics within the LMPC framework. This approach exhibits increased robustness against parameter variations and limited data compared to prior LMPC implementations. While LMPC offers significant advantages, it also has limitations. These include high computational costs due to continuous model updates, dependency on sufficient high-quality training data, and the need for additional hardware for onboard real-time learning and processing. Pure pursuit, another control algorithm used in racing tasks, is simpler and more computationally efficient [217]. It is particularly suitable for scenarios with limited computational resources. However, it lacks the adaptability needed for dynamic environments and may struggle in high-speed racing where precise trajectory prediction is essential.
One of the significant challenges in autonomous racing is localization. Current methods, such as motion capture systems and pre-mapped tracks, rely heavily on external infrastructure [248, 249]. This dependence creates scalability issues and imposes hardware and environmental constraints. To overcome these limitations, future research should focus on developing map-free or vision-based localization systems. Lightweight neural networks could also be leveraged for real-time learning and decision-making, reducing the reliance on extensive infrastructure.